# A Comprehensive Analysis of Social Tie Strength: Definitions, Prediction Methods, and Future Directions

### v2

@@ I II III IV V VI VII
In the above definitions, $\Theta\in(0,\infty)$ is the global threshold, and $\mathbf{W}_{i}(\theta)$ is the top $\theta$ largest edge weight in $[\mathbf{W}_{iv},v\in\mathcal{N}_{i}]$. For each definition that requires a threshold, we first select a range of possible thresholds. For each threshold, we calculate the number of strong ties. Next, we compute the average number of strong ties across all thresholds. Finally, the threshold that produces the number of strong ties closest to this average is chosen as the final threshold. All the final thresholds can be found in the code. Based on these, we can generate the pseudo-ground-truth tie strength labels.
The comparisons among these definitions are shown in Table 1. As the semantic meaning of bidirectionality and weight information can be adjusted easily according to different datasets, we believe these 7 definitions can cover existing definitions of most prior works. Moreover, the evaluations from these definitions can reveal how the models will perform in real-world scenarios.
3.2 Comparison of Definitions
In this study, we collect three real-world social networks. We first introduce some of their details and then we investigate different tie strength definitions on these datasets. Specifically, we compare the strong/weak tie distribution, dive into the relationship between definitions through their correlations, and investigate the resilience of strong and weak ties according to diverse definitions in the Twitter dataset.
Curated Datasets
Table 2: The percentage of ties labeled as strong according to various tie strength definitions for each dataset.
@@ 3.2 Comparison of Definitions
To cover different types of real-world social networks, we selected three datasets: Bitcoin_alpha (Derr et al. 2017) (a business transaction network), CollegeMSG (Panzarasa, Opsahl, and Carley 2009) (a college social network), and Twitter (Kheiri et al. 2023) (an online social network). These datasets range from public to private interactions and contain rich features for our experiments. Below are the dataset details:
Bitcoin_alpha: A cryptocurrency transaction network with anonymous online users as the nodes and their varying levels of trust/distrust ratings as directed weighted edges. We use text embeddings from user comments and filter out distrust edges to focus on trust-based strong/weak ties.
CollegeMSG: A social network of college students, where nodes represent users and edge weights are calculated based on the number of messages sent from one user to another.
Twitter: A following network where nodes are users and edges represent interactions (i.e., retweets and quotes) from one user to another. To ensure we are studying active users, we only include users that have at least posted or deleted a tweet in the prior 3 weeks. We also remove nodes that never interacted with other nodes to ensure the local threshold has practical meaning for them. The edge features are set to be the average embedding of the interacted tweet from source to target nodes.
In each of the resulting networks, we assigned an undirected tie between each pair of nodes with strength determined by the corresponding definitions. The tie weight features were obtained by averaging embeddings between two nodes for the Bitcoin_alpha and Twitter datasets, and by taking the sum of edge weights for CollegeMSG, as no semantic edge features are available for this dataset. This transformation prepares the dataset for the subsequent experiments.
Strong/Weak Tie Distribution
The definitions in Section 3.1 are applied to these datasets to create the pseudo ground-truth tie strength label. The ratios of pseudo ground-truth strong ties under each definition for each dataset are shown in Table 2. We can see that in most cases, these strong ties are the minority group in social networks. This is because strong ties usually serve as inter-community connections, and for most people, their active social circles are likely to be small in comparison with the whole network size. To further verify the effectiveness of our definitions, in Figure 1, we show the average weight, which is calculated as the sum of the weights of all edges between the endpoint nodes of the ties for both strong and weak ties. As shown in Figure 1, the average weight of strong ties is significantly higher than that of weak ties, demonstrating that our definitions successfully encapsulate the difference in strength between strong and weak ties.
Figure 1: Average weights of pseudo ground-truth of strong and weak ties. The average weight of strong ties is significantly higher than the average weight of weak ties.
@@
Figure 2: The Pearson correlation heatmaps (top row) and Euclidean distance heatmaps (bottom row) for Bitcoin_alpha, CollegeMSG, and Twitter under our definitions.
Correlation Between Definitions
We also use the heatmaps of Pearson correlation and Euclidean distance for these three datasets under various definitions to show their relationship. As shown in Figure 2, Definition I is distinct from other definitions as it is the only one that does not consider edge weight as a criterion. Definitions II and III are more similar to each other than to other definitions because they both rely on edge weight on a global scale. Additionally, they require a tie to meet a global threshold, which is set to make sure in most cases only a small fraction of ties can achieve. This makes such ties more likely to be bidirectional. This also stands for Definition VI and VII in our two interaction networks, Bitcoin_alpha and CollegeMSG datasets. However, the Twitter network is a following network based on a large online social network, where one direction followings are common (as shown in Definition I in Table 2). One user’s frequent interaction with another does not necessarily indicate mutual following (e.g. a huge fan of some celebrities, or active users who frequently comment on a topic), but this is much more rarely the case in Bitcoin_alpha and CollegMSG as people who don’t know another are less likely to trade or message frequently. This explains why Definition VI and VII are not so similar in the Twitter network as compared with the Bitcoin_alpha and CollegeMSG networks.
Strong/Weak Tie Resilience
Table 3: The ratio of pseudo ground-truth strong/weak ties in Twitter that became disconnected after $\Delta T$ weeks. R is the average of the ratio between weak and strong ties value.
$\Delta T=4$ $\Delta T=8$ $\Delta T=12$ R
@@
Strong ties are expected to be less likely to break compared to weak ties. To demonstrate that our definitions align with this expectation, we use the Twitter dataset as an example and illustrate the disconnection rates of pseudo ground-truth strong and weak ties after $\Delta T$ weeks in Table 3. Overall, strong ties dissolved less often compared with weak ties across all definitions, and both strong and weak ties have more chances to dissolve with the increase of $\Delta T$. More specifically, Definitions I, III, and VII, which characterize strong ties as bidirectional, exhibit lower disconnection rates, affirming that bidirectional ties are more durable than unidirectional ones. Additionally, as $\Delta T$ increases, the probability of both strong and weak ties breaking rises. According to Figure 3, the ratio of weak to strong tie disconnection rates grows with $\Delta T$, indicating that weak ties break at a faster rate than strong ties. These findings not only corroborate our intuitive understanding but also offer valuable insights that (1) stronger ties are more resilient than weak ties, and (2) this tie resiliency ratio increases as the network evolves. These are important insights into tie strength and we believe that they can pave the way for future advancements in this field.
@@
Figure 3: The disconnected rates ratio between weak over strong ties of each definition after $\Delta T$ weeks. Overall the ratio is becoming larger with the increase of $\Delta T$.
Table 4: Prediction accuracy of existing models on three datasets across seven definitions in the random weakly supervised setting. Overall models perform worse in Definition I, V, and VII, and MLP w/ $\mathbf{E}$ and GCN perform best.
MLP w/o $\mathbf{E}$ 0.653 0.532 0.756 MLP w/o $\mathbf{E}$ 0.521 0.610 0.731
MLP w/ $\mathbf{E}$ 0.789 0.845 0.909 MLP w/ $\mathbf{E}$ 0.560 0.845 0.743
GCN 0.719 0.824 0.885 GCN 0.537 0.809 0.713
GTN 0.640 0.552 0.695 GTN 0.507 0.517 0.624
MLP w/o $\mathbf{E}$ 0.794 0.758 0.930 MLP w/o $\mathbf{E}$ 0.741 0.781 0.573
MLP w/ $\mathbf{E}$ 0.783 1.000 0.939 MLP w/ $\mathbf{E}$ 0.726 0.900 0.544
GCN 0.794 0.976 0.931 GCN 0.674 0.863 0.637
GTN 0.789 0.757 0.931 GTN 0.650 0.794 0.558
MLP w/o $\mathbf{E}$ 0.828 0.789 0.865 MLP w/o $\mathbf{E}$ 0.766 0.830 0.865
MLP w/ $\mathbf{E}$ 0.816 0.965 0.867 MLP w/ $\mathbf{E}$ 0.756 0.907 0.867
GCN 0.828 0.939 0.960 GCN 0.731 0.907 0.867
GTN 0.825 0.789 0.960 GTN 0.666 0.830 0.841
MLP w/o $\mathbf{E}$ 0.889 0.882 0.843 MLP w/o $\mathbf{E}$ 0.742 0.740 0.808
MLP w/ $\mathbf{E}$ 0.882 0.890 0.845 MLP w/ $\mathbf{E}$ 0.759 0.907 0.830
GCN 0.889 0.905 0.845 GCN 0.739 0.889 0.834
GTN 0.886 0.894 0.845 GTN 0.709 0.733 0.779
4 Tie Strength Prediction Methods
In this section, we introduce the real-world settings and models that are used for later experiments.
4.1 Experiment Settings
Based on the definitions in Section 3.1, we can now assign the pseudo ground-truth tie strength in networks. However, as illustrated in Section 1, the ground-truth tie strength information is usually missing or incomplete in reality, and most online platforms lack mechanisms to distinguish between different types of ties. Therefore, when evaluating the existing tie strength prediction models’ performance, the common supervised learning setting with many labeled edges may not align well.
@@ 4.1 Experiment Settings
A more realistic setting would be weakly supervised settings, which stands for the ground-truth label is incomplete, or even unsupervised learning, where all ground-truth information is missing. While the traditional weakly supervised settings is to randomly select training edges from the overall network, a node-centric training selection might better align with the real-world social science domain; i.e., the idea of conducting surveys from a few individuals that provide their tie strengths (compared to randomly distributed user pairs).
Based on these thoughts, we proposed three experiment settings for this study:
@@
Table 5: Prediction accuracy of existing models on three datasets across seven definitions in node-centric weakly supervised setting. Similarly, overall models perform worse in Definition I, V, and VII, and MLP w/ $\mathbf{E}$ and GCN perform best.
MLP w/o $\mathbf{E}$ 0.746 0.506 0.663 MLP w/o $\mathbf{E}$ 0.526 0.638 0.735
MLP w/ $\mathbf{E}$ 0.833 0.817 0.914 MLP w/ $\mathbf{E}$ 0.533 0.813 0.736
GCN 0.741 0.817 0.757 GCN 0.516 0.798 0.733
GTN 0.650 0.511 0.743 GTN 0.495 0.588 0.685
MLP w/o $\mathbf{E}$ 0.794 0.757 0.931 MLP w/o $\mathbf{E}$ 0.741 0.794 0.454
MLP w/ $\mathbf{E}$ 0.794 0.918 0.931 MLP w/ $\mathbf{E}$ 0.729 0.877 0.458
GCN 0.794 0.884 0.931 GCN 0.741 0.855 0.614
GTN 0.760 0.758 0.931 GTN 0.712 0.776 0.524
MLP w/o $\mathbf{E}$ 0.828 0.789 0.960 MLP w/o $\mathbf{E}$ 0.766 0.829 0.867
MLP w/ $\mathbf{E}$ 0.828 0.969 0.960 MLP w/ $\mathbf{E}$ 0.766 0.883 0.867
GCN 0.828 0.885 0.960 GCN 0.766 0.832 0.853
GTN 0.805 0.789 0.960 GTN 0.736 0.830 0.853
MLP w/o $\mathbf{E}$ 0.889 0.894 0.845 MLP w/o $\mathbf{E}$ 0.756 0.744 0.779
MLP w/ $\mathbf{E}$ 0.889 0.894 0.845 MLP w/ $\mathbf{E}$ 0.767 0.882 0.816
GCN 0.889 0.894 0.845 GCN 0.754 0.852 0.813
GTN 0.882 0.894 0.803 GTN 0.720 0.735 0.786
@@
Note that for simplicity, the two weakly supervised settings are all transductive, and all edges in the training set are considered labeled. Inspired by the N-way K-shot setting in few-shot learning (Cheng et al. 2019), we explicitly set the size of the validation set to be equal (approximately equal in node-centric weakly supervised setting) to the size of the training set. To make sure the results between weakly supervised settings I and II are comparable, we explicitly set the number of training edges of these two settings to be approximately the same. For random weakly supervised setting, the training edge set sizes are 20 for Bitcoin_alpha, 70 for CollegeMSG, and 50 for Twitter. For node-centric weakly supervised setting, the number of hub nodes is set to 5 for Bitcoin_alpha, 4 for CollegeMSG, and 2 for Twitter, resulting in an average of 21, 68.3, and 52.3 training edges, respectively, across three splits.
4.2 Models for Tie Strength Prediction
@@
As shown in Section 2.1, this study considers three main approaches to infer tie strength: heuristic methods, STC-based methods, and neural network methods. We focus on using structural features and edge features. For heuristic methods, we adopted the approach from  (Mattie et al. 2018) by training a Random Forest model (Breiman 2001) using structural features such as common neighbors, sum of end nodes’ degrees, sum of clustering coefficients, etc. For STC-based methods, we used STC-Greedy code from (Adriaens et al. 2018). For neural network methods, inspired by  (Zhuang et al. 2021), we selected two variants of a 3-layer Multi-layer Perception (MLP) (Taud and Mas 2018), one that includes edge features and one that does not, and two Graph Neural Networks (GNNs): GCN (Kipf and Welling 2017) and GTN (Shi et al. 2020), where one key difference is whether edge features are utilized in the message-passing process.
5 Empirical Study of Tie Strength Prediction
@@
In this section, we present the results of existing models and offer an in-depth discussion. Following the setting of  (Gilbert and Karahalios 2009; Sohrabi and Akbari 2016), we adopted accuracy as the evaluation metric. Additionally, to reflect the intensity of strong and weak ties, similarly to (Sintos and Tsaparas 2014; Adriaens et al. 2018), we showed the average tie weight differences between strong and weak ties. All results are averaged over 3 runs with different data splits. The experiments were performed on a desktop with a 24GB NVIDIA GeForce RTX 3090 GPU and 64GB RAM. Note that we explicitly show the ”Avg” below each definition to reflect the average performance of the applied method, and use the last “Avg” section in Table 4 and  5 to reflect the average performance of each method on each dataset. Code and datasets are provided at https://github.com/XueqiC/tie˙strength˙prediction.
5.1 Random Weakly Supervised Setting

### v5

@@ I II III IV V VI VII
In these definitions $\Theta\in(0,\infty)$ is the global threshold, and $\mathbf{W}_{i}(\theta)$ is the top $\theta$ largest edge weight in $[\mathbf{W}_{iv},v\in\mathcal{N}_{i}]$. For each definition that requires a threshold, we first select a range of possible thresholds. For each threshold, we calculate the number of strong ties. Next, we compute the average number of strong ties across all thresholds. Finally, the threshold that produces the number of strong ties closest to this average is chosen as the final threshold. All the final thresholds can be found in the code. Based on these, we can generate the pseudo tie strength labels.
The comparisons among these pseudo-labels are shown in Table 1. As the semantic meaning of bi-directionality and weight information can be adjusted easily according to different datasets, we believe these 7 definitions can cover existing definitions in prior works. Moreover, the evaluations from these definitions can reveal how the models will perform in real-world scenarios.
4. Tie Strength Dataset Collection
In this study, we collect three real-world networks. We first introduce some of their details and then conduct an initial analysis. Specifically, we compare their strong/weak tie distribution and then dive into the relationship between TS definitions.
4.1. Dataset Curation
To cover different types of real-world social networks, we selected three datasets, including an anonymous transaction network, a college messaging network, and a popular online social network. These datasets range from public to private interactions and contain rich features for our experiments. Below are the dataset details:
Bitcoin_alpha (Derr et al., 2017; Arya et al., 2025): A cryptocurrency transaction network with anonymous online users as the nodes and their varying levels of explicitly defined trust/distrust ratings as directed weighted edges. We use text embeddings from user comments and filter out distrust edges to focus on trust-based strong/weak ties.
CollegeMSG (Panzarasa et al., 2009): This dataset is a college student interaction social network, where nodes represent students, a directed edge represents a user interacting with another user, and the edge weights represent the number of messages sent from one user to another.
Twitter (Kheiri et al., 2023; Farokhi et al., 2024): This dataset represents a following network, where nodes are users and edges represent interactions between them (e.g., retweets and quotes). During the data collection process, we first identified users using a breadth-first search starting from a selected Twitter account, initially collecting data for 130,000 users, which includes social ties, user-generated content (such as tweets, retweets, and quotes), and, when available, user self-declared bios. To focus on active users, we only include those who have posted or deleted at least one tweet within the past three weeks. Additionally, we remove users who never interacted with others (i.e., no retweets or comments), ensuring that all remaining nodes have meaningful network connections. Lastly, we use the average embedding of tweets that the source user directed to the target user as the edge features, which captures the semantic context of their interaction.
Table 2. The percentage of ties labeled as strong according to various tie strength definitions for each dataset.
@@ 4.1. Dataset Curation
In each of the resulting networks, we assigned an undirected tie between each pair of nodes with strength determined by the corresponding definitions. The tie weight features were obtained by averaging embeddings between two nodes for the Bitcoin_alpha and Twitter datasets, and by taking the sum of edge weights for CollegeMSG, as no semantic edge features are available for this dataset. This transformation prepares the dataset for the experiments.
4.2. Dataset Analysis
In this section, we apply the definitions in Section 3 to the curated datasets to create the pseudo ground-truth tie strength labels. Then we present the dataset analysis to show insights related to label distribution and correlations between definitions.
4.2.1. Label Distribution
The ratios of pseudo ground-truth strong ties under each definition for each dataset are shown in Table 2. We can see that in most cases, these strong ties are the minority group in social networks. This is because strong ties usually serve as inter-community connections, and for most people, their active social circles are likely to be small in comparison with the whole network size. To further verify the effectiveness of our definitions, in Figure 1, we show the average weight, which is calculated as the sum of the weights of all edges between the endpoint nodes of the ties for both strong and weak ties. As shown in Figure 1, the average weight of strong ties is significantly higher than that of weak ties, demonstrating that our definitions successfully encapsulate the difference in strength between strong and weak ties.
Figure 1. Average weights of pseudo ground-truth of strong and weak ties. The average weight of strong ties is significantly higher than the average weight of weak ties.
@@ 4.2.1. Label Distribution
Figure 2. Tie strength definition relation heatmaps using Pearson correlation (top) and Euclidean distance (bottom).
4.2.2. Correlation Between Definitions
We also use the heatmaps of Pearson correlation and Euclidean distance for these three datasets under various definitions to show their relationship. As shown in Figure 2, Definition I is distinct from other definitions as it is the only one that does not consider edge weight as a criterion. Definitions II and III are more similar to each other than to other definitions because they both rely on edge weight on a global scale. Additionally, they require a tie to meet a global threshold, which is set to make sure only a small fraction of ties can achieve this in most cases. This makes such ties more likely to be bidirectional. This also stands for Definition VI and VII in our two interaction networks, Bitcoin_alpha and CollegeMSG datasets. However, the Twitter network is a following network based on a large online social network, where one direction followings are common (as shown in Definition I in Table 2). One user’s frequent interaction with another does not necessarily indicate mutual following (e.g. a huge fan of some celebrities, or active users who frequently comment on a topic), but this is much more rarely the case in Bitcoin_alpha and CollegMSG as people who don’t know another are less likely to trade or message frequently. This explains why Definition VI and VII are not so similar in the Twitter network as compared with the Bitcoin_alpha and CollegeMSG networks.
5. Pseudo-Label Evaluation Framework
Now that we have the pseudo-label techniques, how can we determine which ones are more appropriate for a new dataset? Inspired by the insights that stronger ties are expected to be less likely to break compared to weak ties (Kivran-Swaine et al., 2011), we develop a novel Pseudo-Label Evaluation Framework to evaluate the quality of assigned pseudo-label from a perspective of tie resilience, where the expectation is that strong ties are more resilient and harder to dissolved (disconnected) than weak ties after time. Here we specifically define the tie resilience ratio $R$ as:
(1) $R=\frac{\text{Weak tie dissolve rate}}{\text{{Strong tie dissolve rate}}}$
Then, the pseudo-label evaluation framework is proposed here:
  1. (1)
    Initial Tie Strength Assignment: At time $t_{0}$, assign pseudo-labels to ties using the proposed method and record the counts of strong and weak ties.
  2. (2)
    Tracking Tie Dissolution: After a time interval $\Delta T$, examine the set of ties originally present at $t_{0}$. Identify how many strong/weak ties have dissolved and calculate the strong/weak tie dissolved rates and the tie resilience ratio.
  3. (3)
    Repeated Observations: Repeat the process for varying $\Delta T$ time intervals and record the results.
Table 3. The ratio of strong/weak ties in our Twitter dataset that became dissolved after $\Delta T$ weeks under each definition. The last column/row shows the average tie resilience ratio.
$\Delta T=4$ $\Delta T=8$ $\Delta T=12$ Avg $R$
@@ 5. Pseudo-Label Evaluation Framework
Ideally, stronger ties should have a lower dissolution rate than weaker ties ($R>1$), and the tie resilience ratio is expected to increase as the network evolves. To examine this, we analyze the Twitter dataset, with results presented in Table 3. Overall, strong ties dissolve less frequently than weak ties across all definitions, and both types of ties become more likely to dissolve as $\Delta T$ increases. More specifically, Definitions I, III, and VII, which define strong ties as bidirectional connections, exhibit lower disconnection rates, reinforcing the idea that bidirectional ties are more stable than unidirectional ones. Additionally, as $\Delta T$ increases, the probability of both strong and weak ties breaking rises. Figure 3 further confirms this trend, showing that $R$ increases with $\Delta T$, indicating that weak ties deteriorate faster than strong ties. These findings support our intuition and validate the correctness of the evaluation framework.
@@
Figure 3. The dissolved rates between weak over strong ties of each definition after $\Delta T$ weeks. Overall, the ratio is becoming larger with the increase of $\Delta T$.
Based on the results in Table 3, we gain key insights into the quality of pseudo-labels assigned under different definitions. For instance, Definition I (Avg. $R=1.31$) provides a more meaningful representation of tie strength than Definition V (Avg. $R=1.07$). This aligns with our understanding of Twitter’s social structure, where Definition I (users following each other) better represents a strong tie compared to Definition V (one user frequently commenting on another), which may simply reflect a highly engaged follower. This analysis demonstrates that our framework effectively evaluates the quality of pseudo-labels based on the dataset’s characteristics.
6. Benchmark for Tie Strength Prediction
Now, after having defined the proposed tie strength pseudo-labeling techniques, curated dataset collection, and pseudo-label evaluation framework, in this section, we introduce the details of our benchmark for tie strength prediction. Specifically, we present details on multiple real-world TS prediction settings, representative models used for later experiments, and evaluation criteria for the tie strength prediction benchmarking.
6.1. Experiment Settings
@@
Based on Section 3, we can now assign pseudo ground-truth tie strength labels in networks. However, as discussed in Section 1, the ground-truth tie strength information is usually incomplete or missing, and most online platforms lack mechanisms to distinguish between different types of ties. Therefore, when evaluating the existing tie strength prediction models’ performance, the common supervised learning setting with many labeled edges may not align well. A more realistic setting would be weakly supervised, which stands for the ground-truth label is incomplete, or even unsupervised learning, where all ground-truth information is missing. Moreover, compared with the traditional random weakly supervised setting, a node-centric training selection might better align with the real-world social science domain; i.e., the idea of conducting surveys from a few individuals who provide their tie strengths (compared to randomly distributed user pairs). Based on these thoughts, we propose two experimental settings:
@@
In this study, the weakly supervised setting is transductive, and all edges in the training set are considered labeled. Inspired by the N-way K-shot setting in few-shot learning (Cheng et al., 2019), we explicitly set the size of the validation set to be equal to the size of the training set. To make sure the results between weakly supervised settings I and II are comparable, we explicitly set the number of training edges of these two settings to be approximately the same. Specifically, for the random weakly supervised setting, the training edge set sizes are 20 for Bitcoin_alpha, 70 for CollegeMSG, and 50 for Twitter. For the node-centric weakly supervised setting, the number of hub nodes is set to 5 for Bitcoin_alpha, 4 for CollegeMSG, and 2 for Twitter, resulting in an average of 21, 68.3, and 52.3 training edges, respectively, across three splits.
6.2. Models for Tie Strength Prediction
As in Section 2.1, this study considers three main approaches to infer tie strength: heuristic methods, STC-based methods, and neural network methods. We focus on using structural and edge features.
@@ 6.2. Models for Tie Strength Prediction
For heuristic methods, we adopted the approach from  (Mattie et al., 2018) by training a Random Forest model (Breiman, 2001) using structural features such as common neighbors, the sum of end nodes’ degrees, the sum of clustering coefficients, etc. For STC-based methods, we used STC-Greedy code from (Adriaens et al., 2018). For neural network methods, inspired by  (Zhuang et al., 2021), we selected two variants of a 3-layer Multi-layer Perception (MLP) (Taud and Mas, 2018), one that includes edge features and one that does not, and two Graph Neural Networks (GNNs): GCN (Kipf and Welling, 2017) and GTN (Shi et al., 2020), where one key difference is whether edge features are utilized in the message-passing process.
6.3. Evaluation Criteria
Following the evaluation setting of (Gilbert and Karahalios, 2009; Sohrabi and Akbari, 2016), we adopt accuracy as the primary evaluation metric. Additionally, to better quantify the disparity in intensity between strong and weak ties, similar to (Sintos and Tsaparas, 2014; Adriaens et al., 2018), we also use the average tie weight difference as an additional criterion to evaluate the prediction quality.
7. Empirical Study of Tie Strength Prediction
@@
In this section, we present the results of existing models and offer an in-depth discussion. All results are averaged over 3 runs with different data splits. The experiments were performed on a desktop with a 24GB NVIDIA GeForce RTX 3090 GPU and 64GB RAM. Our code and datasets are all publicly available [^1] https://github.com/XueqiC/Awesome-Tie-Strength-Prediction.
7.1. Random Weakly Supervised Setting